ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 151, 2006

Bartal, Y., N. Linial, M. Mendel and A. Naor Limitations to Fréchet’s metric embedding method

111

Besser, A. A short proof of de Shalit’s cup product formula  53
Choi, Y. Homology of the classifying space of Sp(n) gauge groups  167

Elekes, M. and J. Steprāns Chains of Baire class 1 functions and various notions of special trees

179

Ginzburg, D. Certain conjectures relating unipotent orbits to automorphic representations

323

Huang, W. and X. Ye A local variational relation and applications 237

Hubert, P. and S. Lelièvre Prime arithmetic Teichmüller discs

in H(2) 281

Lelièvre, S. See P. Hubert

Linial, N. See Y. Bartal

Mendel, M. See Y. Bartal

Naor, A. See Y. Bartal

Regev, A. and Y. Roichman Statistics on wreath products and generalized binomial-Stirling numbers

189

Roichman, Y. See A. Regev

Rosłanowski, A. and S. Shelah Measured creatures  61

Shelah, S. See A. Rosłanowski

Shustin, E. and I. Tyomkin Patchworking singular algebraic curves I

125

Shustin, E. and I. Tyomkin Patchworking singular algebraic curves II

145

Steprāns, J. See M. Elekes

Strüngmann, L. Baer cotorsion pairs  29
Su, Y. and B. Xin Classification of quasifinite W_{∞}-modules  223
Tabachnikov, S. Tire track geometry: Variations on a theme  1

Tyomkin, I. See E. Shustin

Xin, B. See Y. Su

Ye, X. See W. Huang

Zieschang, P.-H. The exchange condition for association schemes  357